PYQs

Mathematics 3 Probability And Statistics

3 papers · 103403

2025103403

B.Tech 4th Semester Examination, 2024 (Held in Feb 2025)

3 hours70 marks

Instructions

  • The marks are indicated in the right-hand margin.
  • There are NINE questions in this paper.
  • Attempt FIVE questions in all.
  • Question No. 1 is compulsory.

Questions

Ask AI
  1. Q1

    Choose the correct answer of the following (any seven question only):

    a)

    In a bulb factory, machines A, B and C manufactures 60%, 30% and 10% bulbs respectively. 1%, 2% and 3% of the bulbs produced respectively by A, B and C are found to be defective. A bulb is picked up at random from the total production and found to be defective. Find the probability that this bulb was produced by the machine A. (i) 1/31/3 (ii) 2/52/5 (iii) 1/51/5 (iv) 3/43/4

    b)

    A random variable XX has the following probability distribution:

    XX -2 -1 0 1 2 3
    P(X)P(X) 0.1 kk 0.2 2k2k 0.3 kk

    The value of kk is: (i) .001 (ii) .05 (iii) 0.1 (iv) 0.02

    c)

    Which of the following is not a measure of Dispersion? (i) Mean Deviation (ii) Standard Deviation (iii) Quartile Deviation (iv) Average Deviation from mean.

    d)

    The value of variance of a discrete random variable XX is given by: (i) V(X)=E(X2)−(E(X))2V(X) = E(X^2) - (E(X))^2 (ii) V(X)=∫x2f(x)dxV(X) = \int x^2 f(x) dx (iii) V(X)=P(X=x)V(X) = P(X=x) (iv) V(X)=∑x3P(X=x)V(X) = \sum x^3 P(X=x)

    e)

    For a Negative skewed frequency distribution curve, the third central moment: (i) μ3>0\mu_3 > 0 (ii) μ3<0\mu_3 < 0 (iii) μ3=0\mu_3 = 0 (iv) μ3\mu_3 does not exist.

    f)

    For a Symmetrical distribution, the coefficient of Skewness: (i) α3=0\alpha_3 = 0 (ii) α3=3\alpha_3 = 3 (iii) α3=0\alpha_3 = 0 (iv) α3=−1\alpha_3 = -1

    g)

    If XX and YY are two variates, these can be at most: (i) One regression line (ii) Two regression lines (iii) Three regression lines (iv) an infinite number of regression lines

    h)

    If XX follows a binomial distribution with parameters n=20n=20 and p=0.4p=0.4, what is the expected value E(X)E(X)? (i) 8 (ii) 10 (iii) 12 (iv) 6

    i)

    Probability mass function for a binomial distribution with usual notations is: (i) (nk)pk−1q\binom{n}{k} p^{k-1} q (ii) (nk)pkq\binom{n}{k} p^k q (iii) (nk)pkqn−k\binom{n}{k} p^k q^{n-k} (iv) (nk)pkqn−k\binom{n}{k} p^k q^{n-k}

    j)

    The T-test is used to compare: (i) The means of two groups (ii) The medians of two groups (iii) The variances of two groups (iv) The standard deviations of two groups.

    14 marks
    Ask AI
  2. Q2

    Answer the following:

    a)

    Two cards are drawn without replacement from a pack of 52 cards. Find the probability that: (i) both are red (ii) the first is a king and the second is an ace.

    7 marks
    b)

    In a bulb factory, machines A, B and C manufactures 40%, 40% and 20% bulbs respectively. 4%, 2% and 3% of the bulbs produced respectively by A, B and C are found to be defective. A bulb is picked up at random from the total production and found to be defective. Find the probability that this bulb was produced by the machine A.

    7 marks
    14 marks
    Ask AI
  3. Q3

    Answer the following:

    a)

    Find the Expectation of number of heads in three tosses of fair coin. Also find (i) Mean (ii) E(X2)E(X^2) (iii) Variance.

    7 marks
    b)

    The diameter XX of an electric cable is assumed to be continuous random variable with probability density function is: f(x)=6x(1−x),0≤x≤1f(x) = 6x(1-x), 0 \le x \le 1. (i) Check that above is probability density function. (ii) Compute P(0.5≤x≤0.75)P(0.5 \le x \le 0.75).

    7 marks
    14 marks
    Ask AI
  4. Q4

    Answer the following:

    a)

    In a normal distribution 31% of the items are under 45 and 8% are over 64. Find the mean and variance of the distribution. [Given $f(1.4) = 0.42, f(0.5) = 0.19$].

    7 marks
    b)

    The length of telephone conversion is an exponential variate with mean 3 minutes. Find the probability that call (a) end less than 3 min. (b) takes between 3 to 5 min.

    7 marks
    14 marks
    Ask AI
  5. Q5

    Answer the following:

    a)

    Find the value of kk if f(x,y)=k(1−x)(1−y)f(x, y) = k(1-x)(1-y) for 0<x,y<10 < x, y < 1 is to be joint density function.

    7 marks
    b)

    If the joint p.d.f. of (X,Y)(X, Y) is f(x,y)=6e−2x−3y,x≥0,y≥0f(x, y) = 6e^{-2x-3y}, x \ge 0, y \ge 0, find the conditional density of YY given XX.

    7 marks
    14 marks
    Ask AI
  6. Q6

    Answer the following:

    a)

    Fit a second degree parabola to the following data:

    xx 0 1 2 3 4
    yy 1 4 10 17 30
    4 marks
    b)

    A manufacturer claims that only 4% of his products supplied by him are defective. A random sample of 600 products contains 36 defectives. Test the claim of the manufacturer at 5% level.

    10 marks
    14 marks
    Ask AI
  7. Q7

    Samples of sizes 10 and 14 were taken from two normal population of S.D. 3.5 and 5.2. The sample means were found to be 20.3 and 18.6. Test whether the means of the population are same at 5% level of significance. (Given, tt at 5% level of significance is 2.07).

    14 marks
    Ask AI
  8. Q8

    The number of yeast cells counted in a haemocytometer is compared to the theoretical value is given below. Does the experimental result support the theory? χ2(6−1=5 d.f. at 5%)=11.070\chi^2 (6-1 = 5 \text{ d.f. at } 5\%) = 11.070.

    No. of Yeast cells in the square Observed Frequency Expected Frequency
    0 103 106
    1 143 141
    2 98 93
    3 42 41
    4 8 14
    5 6 5
    14 marks
    Ask AI
  9. Q9

    Write short notes on any two of the following:

    a)

    marginal probability distribution

    b)

    Z-Test for significance of difference of two means in case of Large samples.

    c)

    Chi-square test for goodness of fit.

    d)

    Large sample test for Difference of standard deviation.

    14 marks
    Ask AI

2024103403

B.Tech Examination, 2024

3 hours70 marks

Instructions

  • The marks are indicated in the right-hand margin.
  • There are NINE questions in this paper.
  • Attempt FIVE questions in all.
  • Question No. 1 is compulsory.

Questions

Ask AI
  1. Q1

    Choose the correct answer of the following (any seven question only):

    a)

    In a bulb factory, machines A, B and C manufactures 60%, 30% and 10% bulbs respectively. 1%, 2% and 3% of the bulbs produced respectively by A, B and C are found to be defective. A bulb is picked up at random from the total production and found to be defective. Find the probability that this bulb was produced by the machine A. (i) 1/31/3 (ii) 2/52/5 (iii) 1/51/5 (iv) 3/43/4

    b)

    A random variable XX has the following probability distribution:

    XX -2 -1 0 1 2 3
    P(X)P(X) 0.1 kk 0.2 2k2k 0.3 kk

    The value of kk is: (i) .001 (ii) .05 (iii) 0.1 (iv) 0.02

    c)

    Which of the following is not a measure of Dispersion? (i) Mean Deviation (ii) Standard Deviation (iii) Quartile Deviation (iv) Average Deviation from mean.

    d)

    The value of variance of a discrete random variable XX is given by: (i) V(X)=E(X2)−(E(X))2V(X) = E(X^2) - (E(X))^2 (ii) V(X)=∫x2f(x)dxV(X) = \int x^2 f(x) dx (iii) V(X)=P(X=x)V(X) = P(X=x) (iv) V(X)=∑x3P(X=x)V(X) = \sum x^3 P(X=x)

    e)

    For a Negative skewed frequency distribution curve, the third central moment: (i) μ3>0\mu_3 > 0 (ii) μ3<0\mu_3 < 0 (iii) μ3=0\mu_3 = 0 (iv) μ3\mu_3 does not exist.

    f)

    For a Symmetrical distribution, the coefficient of Skewness: (i) a3=1a_3 = 1 (ii) a3=3a_3 = 3 (iii) a3=0a_3 = 0 (iv) a3=−1a_3 = -1

    g)

    If XX and YY are two variates, these can be at most: (i) One regression line (ii) Two regression lines (iii) Three regression lines (iv) an infinite number of regression lines

    h)

    If XX follows a binomial distribution with parameters n=20n=20 and p=0.4p=0.4, what is the expected value E(X)E(X)? (i) 8 (ii) 10 (iii) 12 (iv) 6

    i)

    Probability mass function for a binomial distribution with usual notations is: (i) (nx)pk−1q\binom{n}{x} p^{k-1} q (ii) (nx)pkq\binom{n}{x} p^k q (iii) (nx)pkqn−k\binom{n}{x} p^k q^{n-k} (iv) (nx)pxqn\binom{n}{x} p^x q^n

    j)

    The T-test is used to compare: (i) The means of two groups (ii) The medians of two groups (iii) The variances of two groups (iv) The standard deviations of two groups.

    Ask AI
  2. Q2

    Answer the following:

    a)

    Two cards are drawn without replacement from a pack of 52 cards. Find the probability that: (i) both are red (ii) the first is a king and the second is an ace.

    b)

    In a bulb factory, machines A, B and C manufactures 40%, 40% and 20% bulbs respectively. 4%, 2% and 3% of the bulbs produced respectively by A, B and C are found to be defective. A bulb is picked up at random from the total production and found to be defective. Find the probability that this bulb was produced by the machine A.

    Ask AI
  3. Q3

    Answer the following:

    a)

    Find the Expectation of number of heads in three tosses of fair coin. Also find (i) Mean (ii) E(X2)E(X^2) (iii) Variance.

    b)

    The diameter XX of an electric cable is assumed to be continuous random variable with probability density function is: f(x)=6x(1−x),0≤x≤1f(x) = 6x(1-x), 0 \le x \le 1. (i) Check that above is probability density function. (ii) Compute P(0.5≤x≤0.75)P(0.5 \le x \le 0.75).

    Ask AI
  4. Q4

    Answer the following:

    a)

    In a normal distribution 31% of the items are under 45 and 8% are over 64. Find the mean and variance of the distribution. [Given $f(1.4) = 0.42, f(0.5) = 0.19$].

    b)

    The length of telephone conversion is an exponential variate with mean 3 minutes. Find the probability that call (a) end less than 3 min. (b) takes between 3 to 5 min.

    Ask AI
  5. Q5

    Answer the following:

    a)

    Find the value of kk if f(x,y)=k(1−x)(1−y)f(x, y) = k(1-x)(1-y) for 0<x,y<10 < x, y < 1 is to be joint density function.

    b)

    If the joint p.d.f. of (X,Y)(X, Y) is f(x,y)=6e−2x−3y,x≥0,y≥0f(x, y) = 6e^{-2x-3y}, x \ge 0, y \ge 0, find the conditional density of YY given XX.

    Ask AI
  6. Q6

    Answer the following:

    a)

    Fit a second degree parabola to the following data:

    xx 0 1 2 3 4
    yy 1 4 10 17 30
    4 marks
    b)

    A manufacturer claims that only 4% of his products supplied by him are defective. A random sample of 600 products contains 36 defectives. Test the claim of the manufacturer at 5% level.

    10 marks
    14 marks
    Ask AI
  7. Q7

    Samples of sizes 10 and 14 were taken from two normal population of S.D. 3.5 and 5.2. The sample means were found to be 20.3 and 18.6. Test whether the means of the population are same at 5% level of significance. (Given, tt at 5% level of significance is 2.07).

    14 marks
    Ask AI
  8. Q8

    The number of yeast cells counted in a haemocytometer is compared to the theoretical value is given below. Does the experimental result support the theory? χ2(6−1=5 d.f. at 5%)=11.070\chi^2 (6-1 = 5 \text{ d.f. at } 5\%) = 11.070.

    No. of Yeast cells in the square Observed Frequency Expected Frequency
    0 103 106
    1 143 141
    2 98 93
    3 42 41
    4 8 14
    5 6 5
    14 marks
    Ask AI
  9. Q9

    Write short notes on any two of the following:

    a)

    marginal probability distribution

    b)

    Z-Test for significance of difference of two means in case of Large samples.

    c)

    Chi-square test for goodness of fit.

    d)

    Large sample test for Difference of standard deviation.

    Ask AI

2022103403

B.Tech Examination, 2022

3 hours70 marks

Instructions

  • The marks are indicated in the right-hand margin.
  • There are NINE questions in this paper.
  • Attempt FIVE questions in all.
  • Question No. 1 is compulsory.

Questions

Ask AI
  1. Q1

    Answer any seven of the following questions:

    a)

    Define mutually exclusive events.

    b)

    Prove that P(Aˉ)=1−P(A)P(\bar{A}) = 1 - P(A), where Aˉ\bar{A} is a complementary event of AA.

    c)

    If P(A)=p1,P(B)=p2P(A) = p_1, P(B) = p_2 and P(A∩B)=p3P(A \cap B) = p_3 then find the value of P(Aˉ∪Bˉ)P(\bar{A} \cup \bar{B}) in terms of p1,p2p_1, p_2 and p3p_3.

    d)

    An urn contains 6 white, 4 red and 9 black balls. If 3 balls are drawn at random, find the probability that one is of each colour.

    e)

    A random variable XX follows Binomial distribution with parameters n=40n=40 and p=14p = \frac{1}{4}. Find the mean of XX.

    f)

    Mean of 100 observations is 50 and standard deviation is 10. What will be the new mean and standard deviation, if each observation is multiplied by 3?

    g)

    Four cards are drawn at random from a pack of 52 cards. Find the probability that two are kings and two are queens.

    h)

    Two unbiased dice are thrown. Find the probability that the first dice shows 6.

    i)

    A random variable XX is such that E[(X−1)2]=10E[(X-1)^2] = 10 and E[(X−2)2]=6E[(X-2)^2] = 6. Find the variance of XX.

    j)

    The moment-generating function of a continuous random variable XX be given as MX(t)=(1−t)−9∣∣t∣<1M_X(t) = (1-t)^{-9} \mid |t| < 1. Then calculate its mean and variance.

    Ask AI
  2. Q2

    Answer the following:

    a)

    If two dice are thrown, what is the probability that the sum is (i) greater than 8 and (ii) neither 7 nor 11?

    b)

    For events A1,A2,A3,A4,A5,…,AnA_1, A_2, A_3, A_4, A_5, \dots, A_n, assuming P(⋃i=1nAi)≤∑i=1nP(Ai)P(\bigcup_{i=1}^n A_i) \le \sum_{i=1}^n P(A_i) then prove that- (i) P(\bigcap_{i=1}^n A_i) \ge \sum_{i=1}^n P(A_i) - (n-1)$; (ii) $P(\bigcap_{i=1}^n A_i) \ge 1 - \sum_{i=1}^n P(\bar{A}_i).

    Ask AI
  3. Q3

    A random variable XX has the density function f(x)={cx2,1≤x≤2cx,2<x<30,otherwisef(x) = \begin{cases} cx^2, & 1 \le x \le 2 \\ cx, & 2 < x < 3 \\ 0, & \text{otherwise} \end{cases}. Find- (i) the constant c$; (ii) $P(1 < x < 2)$; (iii) $P(x \ge 3)$; (iv) $P(x < 1)$; (v) $E(2X)$; (vi) $\text{var}(X-3)$; (vii) $E(2X+4).

    Ask AI
  4. Q4

    Answer the following:

    a)

    Find the probability that in tossing a fair coin 3 times there will appear (i) 3 heads, (ii) 2 tails and 1 head, (iii) at least 1 head and (iv) not more than 1 tail.

    b)

    Ten percent of the tools produced in a certain manufacturing process turn out to be defective. Find the probability that in a sample of 10 tools chosen at random exactly two will be defective, by using (i) the Binomial distribution and (ii) Poisson approximation to the Binomial distribution.

    Ask AI
  5. Q5

    A random variable XX has mean 3 and variance 2. Use Chebyshev's inequality to obtain an upper bound for-

    a)

    $P(|X-3| \ge 1)$;

    b)

    P(∣X−3∣≥2)P(|X-3| \ge 2).

    14 marks
    Ask AI
  6. Q6

    The mean lifetime of electric light bulbs produced by a company has in the past been 1120 hours with a standard deviation of 125 hours. A sample of 8 electric light bulbs recently chosen from a supply of newly produced bulbs showed a mean lifetime of 1070 hours. Test the hypothesis μ=1120\mu = 1120 hours against the alternative hypothesis μ<1120\mu < 1120 hours, using a significance level of-

    a)

    0.05;

    b)

    0.01.

    14 marks
    Ask AI
  7. Q7

    It has been found from experience that the mean breaking strength of a particular brand of thread is 275.6 grams with a standard deviation of 39.7 grams. Recently a sample of 36 pieces of thread showed a mean breaking strength of 253.2 grams. Can one conclude at a significance level of (a) 0.05 and (b) 0.01 that the thread has become inferior?

    14 marks
    Ask AI
  8. Q8

    Find the coefficient of correlation and the lines of regression for the following values of xx and yy:

    x:26475x: 2 \quad 6 \quad 4 \quad 7 \quad 5 y:88562y: 8 \quad 8 \quad 5 \quad 6 \quad 2

    14 marks
    Ask AI
  9. Q9

    Given the joint density function of the random variables (X,Y)(X, Y) as f(x,y)={12xe−y;0<x<2,y>00;elsewheref(x, y) = \begin{cases} \frac{1}{2}xe^{-y}; & 0 < x < 2, y > 0 \\ 0; & \text{elsewhere} \end{cases}. Find the probability density function of X+YX+Y. Also find the mean and variance of XX and YY.

    14 marks
    Ask AI

Support BEUConnect

Contributions cover the domain and hosting.

UPI QR code for pritkr@cnrb
pritkr@cnrb

Add to home screen

  1. Tap Share in Safari.
  2. Choose Add to Home Screen.